### v2

# Robustness to Geographic Distribution Shift Using Location Encoders

Abstract
Geographic distribution shift arises when the distribution of locations on Earth in a training dataset is different from what is seen at test time. The most common approaches to tackling geographic distribution shift treat regions delimited by administrative boundaries such as countries or continents as separate domains and apply standard domain adaptation methods, ignoring geographic coordinates that are often available as metadata. This paper proposes the use of location encoders for modeling continuous, learnable domain assignment. We show how both non-parametric sine-cosine encoders and pre-trained location encoders can be used in conjunction with standard domain adaptation methods for improved robustness to geographic distribution shift. Our proposed methods achieve new state-of-the-art results on two geo-tagged remote sensing datasets from the WILDS benchmark. We have made our code publicly available at: https://github.com/crastoru/wilds-geoshift.
1 Introduction
Empirical risk minimization (ERM) assumes a set of independent and identically distributed (i.i.d) training examples. The learned predictor is then expected to generalize well to unseen data at inference time with the assumption that this data is drawn independently from the same distribution. In many real-world applications, these assumptions do not hold [4].
In this paper, we focus on a particular way of violating these assumptions known as subpopulation shift [9]. In this setting, 1) the training data is not i.i.d but rather drawn from many different distributions (domains), and 2) the inference data is drawn from a different mixture of the same domains. Subpopulation shift arises naturally in many remote sensing applications [5], where training and inference data may be drawn from different mixtures of geographic regions. We refer to this setting as geographic distribution shift.
A common approach to learning a predictor that is robust to geographic distribution shift is to treat different regions delimited by administrative boundaries such as countries or continents as different domains and apply standard domain adaptation methods [9]. However, this approach ignores intra-regional diversity and inter-regional similarities, and geographic coordinates that are often available as metadata are discarded.
The central idea of this paper is to use location encoders to model an underlying, continuous domain-assigning function in place of discrete region labels. Location encoders are parametric functions that map geographic coordinates into a learning-friendly higher-dimensional latent space [11]. With the rise of geo-tagged image datasets [17], location encoders have become a widely applicable tool and can be used to encode a rich domain-specific signal. In particular, geographic coordinates have proven to be a useful input for a variety of prediction tasks, including geo-localization [18], image classification [1], regression [8], and super-resolution [12].
In this paper, we show how location encoders can be naturally integrated into various domain adaptation methods for improved robustness to geographic distribution shift. We conduct experiments using two remote sensing datasets from the WILDS benchmark: FMoW [3] and PovertyMap [21]. Empirically, we demonstrate that our proposed methods lead to significant improvements on worst-group performance metrics under geographic distribution shift. Notably, we achieve a new state-of-the-art result on the WILDS leaderboard for both datasets.
2 Background
2.1 Subpopulation Shift
What can go wrong when learning a predictor $f:X\rightarrow Y$ using ERM in the midst of a subpopulation shift? Intuitively, different domains may have different sets of features that are predictive of the target $y$, and the learned predictor must rely on the correct set of features for each domain without becoming dependent on spurious correlations. Standard ERM can fail in this setting by either underfitting to predictive features from under-represented domains in the training data or by overfitting to spurious features from over-represented domains. In both cases, the result is a predictor that performs significantly better on some domains than others at inference time. This bias is problematic when the mixture of domains at inference time is significantly different from what is seen during training or when the bias can have unfair or harmful consequences when deployed [15].
Many approaches to learning a predictor that is robust to subpopulation shift fall into one of three categories. The first seeks to strengthen the reliance of the predictor on domain-specific predictive features by using domain labels to condition the hypothesis space. Examples include feature-wise modulation [10] and mixture of experts models [20]. The second seeks to reduce the reliance of the predictor on spurious correlations by learning features that satisfy some invariance condition across domains [16] [2] [6]. Note that some recent approaches are more targeted and can be considered part of both categories [19] [7]. The third category consists of distributionally robust optimization approaches that aim to minimize risk according to an estimated worst-case data distribution [15]. In this work, we focus on a subset of approaches from the first two categories, which we describe below.
2.2 Conditional Modeling Methods
2.2.1 Domain Conditional Predictors (DCP)
The key idea behind domain conditional predictors [10] is to learn a latent representation of each domain and to condition the prediction of $y$ on these domain representations using FiLM [13]. The latent domain representations are obtained from an intermediate layer in a neural network that is trained to predict the domain label directly from the input $x$. This network is trained jointly with the predictor $f$. Please refer to Appendix A for the exact structure of the DCP loss function.
2.2.2 Domain Distances for OOD Generalization (D3G)
Given domain labels at training time, D3G [20] trains separate prediction heads for each domain in the training data, and jointly trains another model to predict pairwise similarities between domains (called domain relations) from metadata. At inference time, the predictions output from various heads are weighted by their relation to the input domain. The domain relation predictor is of the form:
$\frac{1}{R}\sum_{r=1}^{R}S_{C}\left(~w_{r}\cdot g(m_{i}),~~w_{r}\cdot g(m_{\text{head {j}}})~\right)$
where $S_{C}$ denotes cosine similarity, $m_{i}$, $m_{\text{head {j}}}$ is metadata representing domain $i$ and the domain of the head $j$ respectively, $w_{r}$} is a set of learnable vectors, and $g$ is a neural network. When applied to geographic distribution shift, D3G uses continent-level labels as the domain metadata input to $g$.
2.3 Invariance-Based Methods
Invariance-based domain adaptation methods seek to learn features that satisfy some invariance condition across domains during training. In this work, we focus on two methods that enforce invariance via a regularization term in the training loss:
  • Invariance risk minimization (IRM): The IRM [2] regularizer encourages features such that a simple classifier common across domains on top of those features may be optimal for all domains in the training data.
  • Correlation Alignment (CORAL): The CORAL [16] regularizer encourages domain-pairwise differences between feature means and covariances to be small.
Figure 1: Illustration of our proposed conditional modeling approaches, location-aware DCP (left) and location-aware D3G (right). The domain predictor in dashed red outline is discarded at inference time.
3 Methods
In this section, we explain how we propose to use location encoders to improve the standard domain adaptation methods described in the previous section for the case of geographic distribution shift.
3.1 Location-Aware Conditional Modeling
3.1.1 Location-Aware DCP
We hypothesize that in the case of geographic distribution shift, geographic coordinates are better predictors of domain than inputs $x$. We therefore propose to substitute the domain predictor in the original DCP method with 2 linear layers trained on top of a simple sine-cosine encoder which we refer to as WRAP, following [1] and [11]:
$\text{WRAP}(\phi,\theta)=\left[~~\sin(\phi)~~\cos(\phi)~~\sin(\theta)~~\cos(\theta)~~\right]$
where $(\phi,\theta)$ is the longitude and latitude. We illustrate this approach in Figure 1. It can be interpreted as a way to condition the predictor $f$ on a location embedding that is encouraged to be predictive of the corresponding domain label. Note that any location encoder features can be used in conjunction with DCP. We choose to experiment with WRAP due to its simplicity, and we compare its performance to that of pre-trained encoders SatCLIP [8] and GeoCLIP [18] whose weights are kept frozen.
3.1.2 Location-Aware D3G
We hypothesize that in the case of geographic distribution shift, geographic coordinates are better predictors of domain relation than discrete region labels. We therefore propose to modify the original D3G method to use geographic coordinates for the metadata $m_{i}$ input to the domain relation predictor, and to use a linear layer trained on top of WRAP features to encode the input domain in place of $g$ (while $m_{\text{head {j}}}$ and its encoding are unchanged). This can be interpreted as a way of predicting domain relations from location embeddings that have been encouraged to align more closely to the corresponding domain label. As with the previous method, we also experiment with frozen pre-trained encoders SatCLIP and GeoCLIP in place of WRAP.
3.2 Location-Aware Invariance-Based Methods
We hypothesize that pre-trained location embeddings that capture the visual characteristics of given locations across the globe can model domain assignments that are more suitable for imposing invariance conditions than administrative boundaries. We propose to use pre-trained location embeddings for clustering locations, and to treat each cluster as a separate domain. Specifically, we perform $k$-means clustering on normalized embeddings for all locations in the training data, where $k$ is a hyperparameter. For validation and inference, we use nearest-neighbor cluster assignment, i.e. we do not include validation and test sets when clustering. We experiment with SatCLIP and GeoCLIP since they are pre-trained on global data.
4 Experiments
DA Method Loc. Encoder Avg. Acc. ($\uparrow$) Worst Acc. ($\uparrow$)
ERM - 67.1 (0.07) 47.6 (0.86)
D3G - 65.8 (0.24) 50.4 (0.92)
IRM - 64.7 (0.12) 48.4 (1.34)
CORAL - 65.3 (0.25) 49.4 (0.73)
ERM WRAP 66.7 (0.34) 47.5 (0.55)
ERM SatCLIP 65.8 (0.98) 50.1 (0.77)
ERM GeoCLIP 66.37 (0.43) 53.4 (1.18)
DCP WRAP 66.9 (0.21) 51.8 (0.86)
DCP SatCLIP 64.6 (0.48) 47.16 (2.15)
DCP GeoCLIP 65.8 (0.68) 51.0 (0.76)
D3G WRAP 66.1 (0.26) 53.8 (0.22)
D3G SatCLIP 63.6 (0.99) 51.5 (1.64)
D3G GeoCLIP 66.5 (0.59) 53.6 (0.06)
IRM SatCLIP 65.2 (0.58) 50.6 (0.76)
IRM GeoCLIP 65.0 (0.72) 49.6 (1.74)
CORAL SatCLIP 65.2 (0.58) 50.6 (0.76)
CORAL GeoCLIP 66.2 (0.97) 50.1 (3.11)
Table 1: Results on the WILDS-FMoW OOD test split. Following the WILDS guidelines, results are averaged across 3 random seeds. Standard error is reported in parentheses.
DA Method Loc. Encoder Avg. r ($\uparrow$) Worst r ($\uparrow$)
ERM - 0.78 (.02) 0.45 (.02)
D3G - 0.76 (.03) 0.37 (.04)
IRM - 0.77 (.02) 0.41 (.04)
CORAL - 0.76 (.03) 0.39 (.05)
ERM WRAP 0.77 (.02) 0.44 (.03)
ERM SatCLIP 0.80 (.02) 0.50 (.04)
ERM GeoCLIP 0.80 (.01) 0.52 (.04)
DCP WRAP 0.78 (.02) 0.51 (.04)
DCP SatCLIP 0.82 (.02) 0.54 (.04)
DCP GeoCLIP 0.82 (.01) 0.57 (.03)
D3G WRAP 0.76 (.02) 0.43 (.03)
D3G SatCLIP 0.78 (.02) 0.48 (.03)
D3G GeoCLIP 0.76 (.03) 0.43 (.06)
IRM SatCLIP 0.76 (.03) 0.41 (.04)
IRM GeoCLIP 0.77 (.03) 0.45 (.03)
CORAL SatCLIP 0.78 (.02) 0.43 (.05)
CORAL GeoCLIP 0.78 (.02) 0.46 (.02)
Table 2: Results on the WILDS-PovertyMap OOD test splits. Following the WILDS guidelines, results are averaged across the 5 official data folds. Standard error is reported in parentheses.
We conduct experiments on two datasets from the WILDS benchmark. FMoW [3] is a land-use classification dataset from satellite imagery covering over 200 countries. Each continent is treated as a different domain. PovertyMap [21] is a dataset of multi-spectral Sentinel-2 imagery for asset wealth prediction, which is a regression task. The data is grouped into two domains corresponding to whether the image was taken in an urban or rural area. For both datasets, the base model we train is an image encoder with a linear prediction head. For FMoW we finetune CLIP ViT-L/14 [14] as the image encoder and for PovertyMap we use a randomly-initialized multi-spectral ResNet-18 as in [9]. A detailed report of hyperparameters can be found in Appendix A.
In Tab. 1 and Tab. 2 we report average and worst-group (domain) performance on the official WILDS OOD test sets for FMoW and PovertyMap respectively. In addition to the standard domain adaptation baselines (D3G, IRM, and CORAL), we experiment with two other baseline methods: standard ERM, and concatenation of location features with the output from the image encoder before the linear prediction head. The results for these baselines are shown in the upper two sections of each table.
We first observe that conditional modeling methods, shown in the third section of each table, appear to be the top-performing category of methods across both datasets. In particular, the top two results for each dataset (D3G for FMoW and DCP for PovertyMap) belong to this category and achieve a new state-of-the-art result on the official WILDS leaderboard. These conditional modeling methods outperform all concatenation baselines, suggesting that location encoders yield the greatest improvement in worst-group performance when used in conjunction with standard domain adaptation methods.
Finally, with a single exception per dataset (DCP + SatCLIP on FMoW and IRM + SatCLIP on PovertyMap), we observe that every proposed location-aware variant of a domain adaptation method outperforms its vanilla counterpart on worst-group performance. Interestingly, all baseline domain adaptation methods applied to PovertyMap lead to a decrease in worst-group performance compared to standard ERM, but most location-aware methods do not. In short, our results provide clear evidence that regardless of the choice of domain adaptation strategy, its location-aware variant should be preferred.
5 Conclusion
It is written in [4] that though sample correlations are ubiquitous in real-world datasets, “we do not appreciate the benefits of modeling these correlations, and the ease with which this can be accomplished algorithmically.” In this paper, we show how location encoders can be leveraged as a simple and effective tool for modeling correlations between the visual characteristics of different regions on Earth. We demonstrate that using them in conjunction with standard domain adaptation methods can greatly improve worst-group performance under geographic distribution shift. Though our methods achieve new state-of-the-art results on two real-world benchmark datasets, we observe that the optimal choices of location features and domain adaptation methods vary between them. Future work may involve evaluating our proposed methods on other tasks beyond regression and classification and understanding the suitability of different location features and domain adaptation methods for different datasets, tasks, and group splits.
Appendix A Appendix
In this section, we describe the training hyperparameters that are specific to each dataset (FMoW and PovertyMap) and specific to each domain adaptation method (DCP, D3G, IRM, and CORAL).
FMoW
Each model is trained for 5 epochs on a single GPU with early stopping on validation loss. We train with batch size 16 and use the Adam optimizer with initial learning rate $10^{-4}$ that decays by a factor of 0.96 each epoch. The CLIP ViT-L/14 backbone is finetuned with the same optimizer and learning rate schedule but with an initial learning rate of $10^{-5}$. We also use gradient accumulation to achieve an effective batch size of 64. Finally, we normalize all images using ImageNet mean and standard deviation, and apply random horizontal flip.
PovertyMap
As in [9], each model is trained for 200 epochs on a single GPU with early stopping on validation $r$. We train with batch size 64 and use the Adam optimizer with initial learning rate $10^{-3}$ that decays by a factor of 0.96 each epoch. As in [9], we apply random horizontal and vertical flip and color jitter to each image.
DCP
The linear layer used for domain prediction is also trained with Adam and step decay learning rate schedule (decayed by a factor of 0.96 per epoch) but with initial learning rate 0.1 times the initial learning rate of the prediction head (i.e. $0.1\times 10^{-4}$ for FMoW and $0.1\times 10^{-3}$ for PovertyMap). The DCP loss function includes two terms: the task prediction (TP) loss, and the domain prediction (DP) loss. It is of the form:
$\mathcal{L}_{TP}+\lambda\mathcal{L}_{DP}$
where $\mathcal{L}$ is classification loss (cross entropy) and $\lambda$ is a hyperparameter. We experimented with values of $\lambda$ in $\{0.,0.1,0.2\}$ and found that $\lambda=0.2$ performed best across all experiments. Note that DCP + WRAP with $\lambda=0$ is equivalent to conditioning prediction on WRAP-encoded location features with FiLM. All pre-trained location encoders used in conjunction with DCP are kept frozen.
D3G
There are two training hyperparameters used in D3G, denoted $\lambda$ and $\beta$ in the original paper. Similar to DCP, D3G has two loss terms: the task prediction loss (denoted $\mathcal{L}_{pred}$ in the original paper) computed from the prediction head for the current domain, and the consistency loss ($\mathcal{L}_{rel}$) that is computed from all other prediction heads whose predictions are weighted by their relation to the current domain. The final loss is of the form:
$\mathcal{L}_{pred}+\lambda\mathcal{L}_{rel}$
The $\beta$ hyperparameter is used to average fixed and learned domain relations (see [20] for more details). We use a value of $\lambda=0.5$ and $\beta=0.8$ for all experiments. All pre-trained location encoders used in conjunction with D3G are kept frozen.
IRM
The IRM loss function includes two terms: the first is the task prediction loss and the second is the IRM penalty term, weighted by $\lambda$ which is a hyperparameter. In all experiments with IRM, we use $\lambda=0.1$ and compute IRM penalty on the final output from the image encoder (before the linear prediction head).
CORAL
The CORAL loss function includes two terms: the first is the task prediction loss and the second is the CORAL penalty, computed as the sum of 1) the squared L2 norms of all domain-pairwise differences between feature means in the current batch, and 2) the squared Frobenius norms of all domain-pairwise differences between feature covariances in the current batch. The second term is weighted by a hyperparameter $\lambda$. In all experiments with CORAL, we use $\lambda=0.1$ and compute the CORAL penalty on the final output from the image encoder (before the linear prediction head).